# Approximations to worst-case data dropping: unmasking failure modes

### v3

@@ 1 Introduction
In the present work, we systematically explore failure modes for approximations to dropping worst-case very-small fractions ($<1\%$) of data. We focus on natural data settings with no adversary. Like [Kuschnig et al., 2021, Moitra and Rohatgi, 2023, Freund and Hopkins, 2023], we focus on linear regression fit with OLS. While linear regression is less common in engineering disciplines, Castro Torres and Akbaritabar [2024] demonstrate that, as recently as 2022, (often well) over half of all papers reporting any methods in Medical and Health Sciences, Agricultural Sciences, Social Sciences, and the Humanities use linear regression. Indeed, most of the applied papers cited in the discussion above use linear regression [Angelucci and De Giorgi, 2009, Finkelstein et al., 2012, Martinez, 2022, Davies et al., 2024, Burton and Roach, 2023]. We suspect OLS is the most common form of linear regression used in practice. In the present work, we identify failure modes in cases of OLS-fit linear regression with one covariate. In light of the the surfaced failure modes, we end with some recommendations for the user and directions for future work.
@@ 3.2 Greedy approximations
Greedy One-Exact. The outlier detection literature has highlighted the combinatorial cost of finding influential subsets exactly. This literature also describes the masking problem that can arise in additive approximations of influence: namely, when one outlier hides the effect of another [Belsley et al., 1980, Atkinson, 1986]. To address these issues, Belsley et al. [1980, Section 2.1] suggest to greedily remove one outlier point at a time in a stepwise procedure. Kuschnig et al. [2021] propose a similar greedy procedure for approximating the Maximum Influence Perturbation; namely, they iteratively: (1) choose one data point to drop using One-Exact, (2) re-run the full analysis without that point. In general, Greedy One-Exact requires $\left\lfloor\alpha N\right\rfloor$ times the cost of an additive approximation. So in the special case of OLS with an effect-size quantity of interest, Greedy One-Exact costs only $\left\lfloor\alpha N\right\rfloor$ data analyses, which is not prohibitive in our examples below. For more general losses though, we expect Greedy One-Exact to cost $\smash{\left\lfloor\alpha N^{2}\right\rfloor}$ times the cost of a single data analysis, which will often be prohibitive, especially as the analyses themselves become more complex and costly.
@@ 4 Failure Modes
Figure 1: Our examples: one-outlier (left), Simpson’s paradox (middle), and poor conditioning (right). A dashed line represents the OLS-estimated slope on the entire dataset while a solid line represents the slope with black dots removed. The tables display the performance of the approximations for the corresponding example: a red X indicates a Type 2 failure, and a green check indicates a success.
@@ 4.1 One outlier example
By definition, Additive One-Exact is exact for removing a single point. However, the AMIP faces Type 1 and 2 failures. There is no distinction between greedy and additive algorithms when $\left\lfloor\alpha N\right\rfloor=1.$ In realistic data settings, we may have a single data point far from the bulk of the data; this outlier may arise due to data-entry errors, machine-measurement errors, or heavy tails in both the covariate and response. To construct the plot in Figure 1 (left), we draw 4,999 red crosses by taking $x_{n}\sim\mathcal{N}(0,1)$ iid and $y_{n}=-x_{n}+\epsilon_{n}$ with $\epsilon_{n}\sim\mathcal{N}(0,1)$ iid. Throughout, we use $\mathcal{N}(\mu,\sigma^{2})$ to denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$. The black dot appears at $x_{n}=y_{n}=100$. The OLS-estimated slope on the full dataset is (nearly) 1; after dropping the black point (0.02% of the dataset), the estimate is (nearly) -1, representing a sign change. AMIP chooses a red cross to drop; it predicts that no sign change will be achieved, and dropping the chosen point and re-running also does not achieve a sign change (Table 1). It follows that AMIP suffers both types of failure in this example.
Intuition. To see what goes wrong, consider that in a single-covariate linear regression with no offset, the influence score is the leverage times the residual divided by the covariate value [Hampel, 1974]. A sufficiently-far outlier will have a very low residual. Meanwhile the leverage is bounded, and dividing by a large covariate will take it toward zero. So a sufficiently-far outlier will have a vanishingly low influence score. See Section D.2 for details. Although previous work has noted that high-leverage observations can lead to problems for algorithms that approximate the Most Influential Set [Kuschnig et al., 2021], they do not demonstrate this result in a setting with one outlier. This one-outlier setting allows us to mathematically analyze why high-leverage points lead to errors in the influence-function approximation. Specifically, we create an instance where the outlier data point goes to infinity in both the x and y direction. In Section D.2, we prove that we can expect such failures when the black-dot point is chosen with $y_{n}=cx_{n}$ for any constant $c>0$ and sufficiently large $|x_{n}|$.
@@ 4.2 Simpson’s paradox
We next see that, if we replace the single outlier in the previous example with a small, noisy group of outliers, both AMIP and Additive One-Exact can suffer Type 1 and 2 failures. The greedy algorithms succeed in this setting. It is common to have (at least) two noisy subpopulations within a single dataset; we consider the case where one subpopulation represents a small fraction of the total. For instance, we might have heterogeneity in the population that the regression model does not account for; such a situation is commonly known as Simpson’s paradox. In the particular example in Figure 1 (center), the overall slope (across all the data) has a different sign than the slope in just the red data or just the black data. To create the illustration in Figure 1 (center), we draw 1,000 red crosses with $x_{n}\sim\mathcal{N}(0,0.25)$ iid, $y_{n}=-x_{n}+\epsilon_{n}$, and $\epsilon_{n}\sim\mathcal{N}(0,1)$ iid. We draw 10 black dots with $x_{n}\sim\mathcal{N}(25,0.25)$, $y_{n}=-x+40+\epsilon_{n}$ iid, and $\epsilon_{n}$ as before. The OLS-estimated slope on the full dataset is 0.52. Dropping the black dots ($0.99\%$ of the data) yields a slope of -0.99, a sign change.
@@ 4.3 Poor conditioning
In Figure 1 (right), we generate the red crosses so as to have poor conditioning; since there is much more noise around the (zero) trend than variation in the covariates, there is no clear regression solution. In particular, we generate the 1,000 red crosses with $x_{n}\sim\mathcal{N}(0,0.001)$ iid, $y_{n}=\epsilon_{n}$, and $\epsilon_{n}\sim\mathcal{N}(0,1)$ iid. We draw the 10 black dots as $x_{n}\sim\mathcal{N}(-1,0.01)$ iid, $y_{n}=-x_{n}-10-\epsilon_{n}$, and $\epsilon_{n}$ as before. When we consider both black dots and red crosses together as a single dataset, there is no poor conditioning. The OLS-estimated slope on the full dataset is around 7.40; dropping the black dots ($0.99\%$ of the data) yields a slope of about -1.04, a sign change.
@@ C.1 AMIP supplementals
Let $e_{p}$ be the $p$th standard basis vector. Then the linear approximation in the setting where the quantity of interest is the sign of the $p$-th regression coefficient becomes
@@ D.2 One-outlier example
Let $(x_{i},y_{i})$ denote the outlier point, where $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$. Let $\|v\|=1$, $e\in\mathbb{R}^{P}$, and $\mathbf{X}\in\mathbb{R}^{NxP}$, and assume that the OLS estimator remains well-defined upon the removal of the outlier (the covariance matrix is non-degenerate). As $\lambda\to\infty$, the influence with respect to $e^{\top}\beta$ goes to zero. Note, this result holds more generally for $y_{i}=c\lambda$, for any constant $c>0$.
@@
$IF((x_{i},y_{i});\hat{\beta})=\underbrace{(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}}_{\text{leverage term}}\underbrace{(y_{i}-\hat{\beta}^{\top}x_{i})}_{\text{residual term}}$ (13)
The leverage term, $(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$, approaches 0 as $\lambda\to\infty$.
$(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$ $=(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i}+x_{i}x_{i}^{\top})x_{i}$
$=((\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}+\frac{(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}})x_{i}$ (Sherman-Morrison formula)
$=\frac{(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}$ (algebraic simplification)
$=\frac{\lambda(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v}{1+\lambda^{2}v^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$)
Hence, the leverage term vanishes at rate $O(\frac{1}{\lambda})$.
@@
$\hat{\beta}^{\top}x_{i}$ $=y^{\top}\mathbf{X}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$ (OLS solution) (14)
$=y^{\top}\mathbf{X}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}-\frac{y^{\top}\mathbf{X}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}$ (Sherman-Morrison formula)
$=y^{\top}\mathbf{X}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}(1-\frac{x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}})$ (algebraic simplification)
$=\frac{y^{\top}\mathbf{X}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}$
$=\frac{(x_{i}y_{i}+\sum_{j\neq i}y_{j}x_{j}^{\top})(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}{1+x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}$
$=\frac{(\lambda^{2}\|v\|^{2}+\sum_{j\neq i}y_{j}x_{j}^{\top})(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}\lambda v}{1+\lambda^{2}v^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$)
$=\frac{\lambda^{2}\|v\|^{2}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v+\lambda\sum_{j\neq i}y_{j}x_{j}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v}{1+\lambda^{2}v^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v}$
Let $A=(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})$.
@@
$\frac{1}{1-x_{i}^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}x_{i}}$ (17)
$=\frac{1}{1-\frac{1}{(\lambda^{2}v^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v)^{-1}+1}}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$)
$=\lambda^{2}v^{\top}(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}v+1.$ (algebraic simplification)
@@
$=((X^{\top}X)^{-1}X^{\top}y-(X_{-i}^{\top}X_{-i})^{-1}X_{-i}^{\top}y_{-i}))$ (Sherman-Morrison formula)
@@
Table 4 shows key regression diagnostics for the One Outlier example presented in Section 4. As the outlier moves far from the other points in both the x and y directions (recall in this example that $x=y$), the residual approaches zero, the leverage, $h_{i}$, approaches one, the leverage correction term $\frac{1}{1-h_{i}}$ approaches infinity, the influence score approaches zero, and the one-step Newton score remains exact for the effect of removing the outlier. Figure 2 shows a leverage plot of the One-Outlier example. The leverage of the extreme outlier is much larger than the leverage of the other data points.
Table 4: Table of Outlier Diagnostics for Extreme Outlier. Each column is defined as follows: Residual: $(y_{i}-\hat{\beta}^{\top}x_{i})$, Leverage: $(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$, Leverage Correction: $\frac{1}{1-x_{i}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}}$ , Influence Score: $(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}(y_{i}-\hat{\beta}^{\top}x_{i})$, One-Step Newton: $\frac{1}{1-x_{i}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}(y_{i}-\hat{\beta}^{\top}x_{i})$.
x Residual Leverage Leverage Correction Influence Score One-Step Newton
10 6.55 0.67 3.00 0.44 1.35
20 4.35 0.89 9.13 0.19 1.77
100 0.98 0.99 204.08 9.78e-3 2.00
1000 0.1 1.00 1.95e4 1.00e-4 2.00
10000 0.0098 1.00 2.04e6 9.83e-7 2.00

### v4

@@ 1 Introduction
In the present work, we systematically explore failure modes for approximations to dropping worst-case very-small fractions (¡ 1%) of data. We focus on natural data settings with no adversary. Like [Kuschnig et al., 2021, Moitra and Rohatgi, 2023, Freund and Hopkins, 2023], we focus on linear regression fit with OLS. Before the present work and contemporaneous work by Hu et al. [2024], there had not been studies systematically characterizing non-adversarial failure modes of approximations to the Most Influential Set and comparing the prominent existing approximations.
In many aspects, Hu et al. [2024] and our present work are complementary. For instance, while Hu et al. [2024] focus on exact recovery of the most influential data subset with size at most equal to a stated value, we focus on finding whether there exists a small subset of data that could change substantive conclusions. Given our different focuses, Hu et al. [2024] finds it useful to separate masking into two phenomena: amplification and cancellation. Meanwhile, we find it useful to point out failure modes due to poor conditioning of the bulk of data. Hu et al. [2024]’s theory assumes a particular data-generating process [their Equation 7]; while we don’t make this assumption, we instead need to take a limit of an outlier data point’s position to derive our results (Proposition D.1).
Both Hu et al. [2024] and our work focus on OLS for theory and illustration of failure modes. While linear regression is less common in engineering disciplines, Castro Torres and Akbaritabar [2024] demonstrate that, as recently as 2022, (often well) over half of all papers reporting any methods in Medical and Health Sciences, Agricultural Sciences, Social Sciences, and the Humanities use linear regression. Indeed, most of the applied papers cited in the discussion above use linear regression [Angelucci and De Giorgi, 2009, Finkelstein et al., 2012, Martinez, 2022, Davies et al., 2024, Burton and Roach, 2023]. We suspect OLS is the most common form of linear regression used in practice. In the present work, we identify failure modes in cases of OLS-fit linear regression with one covariate. In light of the surfaced failure modes, we end with some recommendations for the user and directions for future work.
@@ 3.2 Greedy approximations
Greedy One-Exact. The outlier detection literature has highlighted the combinatorial cost of finding influential subsets exactly. This literature also describes the masking problem that can arise in additive approximations of influence: namely, when one outlier hides the effect of another [Belsley et al., 1980, Atkinson, 1986]. To address these issues, Belsley et al. [1980, Section 2.1] suggest to greedily remove one outlier point at a time in a stepwise procedure. Kuschnig et al. [2021] propose a similar greedy procedure for approximating the Maximum Influence Perturbation; namely, they iteratively: (1) choose one data point that makes the largest exact change when dropped (equivalently, run Additive One-Exact for removing just one data point), (2) re-run the full analysis without that point. In general, Greedy One-Exact requires $\left\lfloor\alpha N\right\rfloor$ times the cost of an additive approximation. So in the special case of OLS with an effect-size quantity of interest, Greedy One-Exact costs only $\left\lfloor\alpha N\right\rfloor$ data analyses, which is not prohibitive in our examples below. For more general losses though, we expect Greedy One-Exact to cost $\smash{\left\lfloor\alpha N^{2}\right\rfloor}$ times the cost of a single data analysis, which will often be prohibitive, especially as the analyses themselves become more complex and costly.
@@ 4 Failure Modes
Figure 1: Our examples: one-outlier (left), Simpson’s paradox (middle), and poor conditioning (right). A dashed line represents the OLS-estimated slope on the entire dataset while a solid line represents the slope with black dots removed. The tables display the performance of the approximations for the corresponding example: a red X indicates a Type 2 failure, and a green check indicates a success. The left plot includes an inset zoomed in on the bulk of the data.
@@ 4.1 One outlier example
By definition, Additive One-Exact is exact for removing a single point. However, the AMIP faces Type 1 and 2 failures. There is no distinction between greedy and additive algorithms when $\left\lfloor\alpha N\right\rfloor=1.$ In realistic data settings, we may have a single data point far from the bulk of the data; this outlier may arise due to data-entry errors, machine-measurement errors, or heavy tails in both the covariate and response. To construct the plot in Figure 1 (left), we draw 1,000 red crosses by taking $x_{n}\sim\mathcal{N}(0,1)$ i.i.d. and $y_{n}=-x_{n}+\epsilon_{n}$ with $\epsilon_{n}\sim\mathcal{N}(0,1)$ i.i.d. Throughout, we use $\mathcal{N}(\mu,\sigma^{2})$ to denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$. The black dot appears at $x_{n}=y_{n}=10^{6}$. We fit OLS with an intercept. The OLS-estimated slope on the full dataset is (nearly) 1; after dropping the black point (0.02% of the dataset), the estimate is (nearly) -1, representing a sign change. AMIP chooses a red cross to drop; it predicts that no sign change will be achieved, and dropping the chosen point and re-running also does not achieve a sign change (Table 1). It follows that AMIP suffers both types of failure in this example.
Intuition. To see what goes wrong, consider that in linear regression, the influence score factorizes into two terms: the residual times a leverage-like term [Hampel, 1974]; see Section D.2 and Equation 13 for full details. A sufficiently-far outlier will have a very low residual. Meanwhile the leverage-like term goes to zero as the outlier gets farther away. So a sufficiently-far outlier will have a vanishingly low influence score. See Section D.2 for details. Although previous work has noted that high-leverage observations can lead to problems for algorithms that approximate the Most Influential Set [Kuschnig et al., 2021], they do not demonstrate this result in a setting with one outlier. This one-outlier setting allows us to mathematically analyze why certain points lead to errors in the influence-function approximation. For instance, our theory suggests that we can achieve similar behavior for outliers in other directions. Specifically, in Section D.2, we consider a point with $y_{n}=cx_{n}$ for any constant $c>0$. We show that, for sufficiently large $|x_{n}|$, the AMIP approximation makes Type 1 errors.
@@ 4.2 Simpson’s paradox
We next see that, if we have a small, noisy group of outliers instead of a single outlier, both AMIP and Additive One-Exact can suffer Type 1 and 2 failures. The greedy algorithms succeed in this setting. It is common to have (at least) two noisy subpopulations within a single dataset; we consider the case where one subpopulation represents a small fraction of the total. For instance, we might have heterogeneity in the population that the regression model does not account for; such a situation is commonly known as Simpson’s paradox. In the particular example in Figure 1 (center), the overall slope (across all the data) has a different sign than the slope in just the red data or just the black data. To create the illustration in Figure 1 (center), we draw 1,000 red crosses with $x_{n}\sim\mathcal{N}(0,0.25)$ iid, $y_{n}=-x_{n}+\epsilon_{n}$, and $\epsilon_{n}\sim\mathcal{N}(0,1)$ iid. We draw 10 black dots with $x_{n}\sim\mathcal{N}(25,0.25)$, $y_{n}=-x_{n}+40+\epsilon_{n}$ iid, and $\epsilon_{n}$ as before. We fit OLS with an intercept. The OLS-estimated slope on the full dataset is 0.52. Dropping the black dots ($0.99\%$ of the data) yields a slope of -0.99, a sign change.
@@ 4.3 Poor conditioning
In Figure 1 (right), we generate the red crosses so as to have poor conditioning; since there is much more noise around the (zero) trend than variation in the covariates, there is no clear regression solution. In particular, we generate the 1,000 red crosses with $x_{n}\sim\mathcal{N}(0,0.001)$ iid, $y_{n}=\epsilon_{n}$, and $\epsilon_{n}\sim\mathcal{N}(0,1)$ iid. We draw the 10 black dots as $x_{n}\sim\mathcal{N}(-1,0.01)$ iid, $y_{n}=-x_{n}-10-\epsilon_{n}$, and $\epsilon_{n}$ as before. We fit OLS with an intercept. When we consider both black dots and red crosses together as a single dataset, there is no poor conditioning. The OLS-estimated slope on the full dataset is around 7.40; dropping the black dots ($0.99\%$ of the data) yields a slope of about -1.04, a sign change.
@@ C.1 AMIP supplementals
Let $e_{p}$ be the $p$th standard basis vector. Then the linear approximation in the setting where the quantity of interest is the sign of the $p$th regression coefficient becomes
@@ D.2 One-outlier example
Let $(x_{i},y_{i})$ denote the outlier point, where $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$. Let $\|v\|=1$, $e\in\mathbb{R}^{P}$, and $\mathbf{X}\in\mathbb{R}^{N\times P}$, and assume that the OLS estimator remains well-defined upon the removal of the outlier (the covariance matrix is non-degenerate). As $\lambda\to\infty$, the influence with respect to $e^{\top}\beta$ goes to zero. Note, this result holds more generally for $y_{i}=c\lambda$, for any constant $c>0$.
@@
$IF((x_{i},y_{i});\hat{\beta})=\underbrace{(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}}_{\text{leverage-like term}}\underbrace{(y_{i}-\hat{\beta}^{\top}x_{i})}_{\text{residual term}}$ (13)
Let $A=(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})$. The leverage-like term, $(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$, approaches $0\in\mathbb{R}^{P}$ as $\lambda\to\infty$.
$(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$ $=(A+x_{i}x_{i}^{\top})^{-1}x_{i}$
$=(A^{-1}-\frac{A^{-1}x_{i}x_{i}^{\top}A^{-1}}{1+x_{i}^{\top}A^{-1}x_{i}})x_{i}$ (Sherman-Morrison formula)
$=\frac{A^{-1}x_{i}}{1+x_{i}^{\top}A^{-1}x_{i}}$ (algebraic simplification)
$=\frac{\lambda A^{-1}v}{1+\lambda^{2}v^{\top}A^{-1}v}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$)
Hence, the leverage-like term vanishes at rate $O(\frac{1}{\lambda})$.
@@
$\hat{\beta}^{\top}x_{i}$ $=y^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$ (OLS solution) (14)
$=y^{\top}\mathbf{X}A^{-1}x_{i}-\frac{y^{\top}\mathbf{X}A^{-1}x_{i}x_{i}^{\top}A^{-1}x_{i}}{1+x_{i}^{\top}A^{-1}x_{i}}$ (Sherman-Morrison formula)
$=y^{\top}\mathbf{X}A^{-1}x_{i}(1-\frac{x_{i}^{\top}A^{-1}x_{i}}{1+x_{i}^{\top}A^{-1}x_{i}})$ (algebraic simplification)
$=\frac{y^{\top}\mathbf{X}A^{-1}x_{i}}{1+x_{i}^{\top}A^{-1}x_{i}}$
$=\frac{(y_{i}x_{i}^{\top}+\sum_{j\neq i}y_{j}x_{j}^{\top})A^{-1}x_{i}}{1+x_{i}^{\top}A^{-1}x_{i}}$
$=\frac{(\lambda^{2}v^{\top}+\sum_{j\neq i}y_{j}x_{j}^{\top})A^{-1}\lambda v}{1+\lambda^{2}v^{\top}A^{-1}v}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$, $y_{i}=\lambda$)
$=\frac{\lambda^{3}v^{\top}A^{-1}v+\lambda\sum_{j\neq i}y_{j}x_{j}^{\top}A^{-1}v}{1+\lambda^{2}v^{\top}A^{-1}v}$
@@
$\frac{1}{1-x_{i}^{\top}A^{-1}x_{i}}$ (17)
$=\frac{1}{1-\frac{1}{(\lambda^{2}v^{\top}A^{-1}v)^{-1}+1}}$ (substituting $x_{i}=\lambda v\in\mathbb{R}^{P}$)
$=\lambda^{2}v^{\top}A^{-1}v+1.$ (algebraic simplification)
@@
$=((\mathbf{X}^{\top}\mathbf{X})^{-1}X^{\top}y-(\mathbf{X}_{-i}^{\top}\mathbf{X}_{-i})^{-1}\mathbf{X}_{-i}^{\top}y_{-i}))$ (Sherman-Morrison formula)
@@
Table 4 and Table 5 display empirical findings for the data generating process described in Section 4.1, showing that a sufficiently far outlier will have vanishingly low influence score (see Proposition D.1). As the black-dot point moves far from the group of red-cross points in both the x and y directions, both the leverage-like term and the residual term of the influence function approach zero at rate $O(\frac{1}{\lambda})$ (see columns 3 and 4 of Table 4). When observing the behavior of the corresponding red-cross point with the largest influence score in Table 5, we see that the leverage-like term approaches zero while the residual term stays relatively constant (within the same order of magnitude). Thus, for sufficiently large values of $(x_{i},y_{i})$, the influence score for the black-dot point becomes smaller than that of a red-cross point (see the highlighted values in Table 4 and Table 5); for the $(x_{i},y_{i})$ values with highlighted influence scores (see Table 4), both AMIP and Greedy AMIP fail (in both the Type 1 and 2 sense). This occurs at $x=y=10^{6}$ for the data generating process described in Section 4.1.
Table 4: This table shows the influence and 1Exact scores for the black-dot point at various values of $(x_{i},y_{i})$ for the data generating process described in Section 4.1 (see plot for the setting where $(x_{i},y_{i})=(10^{6},10^{6})$ in Figure 1 (left)). In order to obtain the influence with respect to $\beta_{1}$, we let $e_{1}=(0,1)$, the standard basis vector corresponding to the x term in our linear regression setup. The Influence Score, $e_{1}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}(y_{i}-\hat{\beta}x_{i})$, is a product between the quantity in column 3 (which we call the leverage-like term, see Equation 13) and column 4 (the residuals). The 1Exact Score is the change in effect size that results from dropping the data point at $(x_{i},y_{i})$ and refitting OLS. The influence scores highlighted in yellow are those that are smaller than the influence score of a red-cross point, leading AMIP and Greedy AMIP to misidentify the Most Influential Set of size 1, resulting in a Type 2 failure.
Black Dot:
$x_{i}$ $y_{i}$ $e_{1}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{i}$ $(y_{i}-\hat{\beta}x_{i})$ Influence Score 1Exact Score
1e1 1e1 9.36e-3 18.390 1.72e-1 1.90e-1
1e2 1e2 9.11e-3 17.981 1.64e-1 1.85
1e4 1e4 1.00e-5 1.97e-1 1.97e-5 2.03
1e6 1e6 1.00e-6 1.97e-3 1.97e-9 2.03
1e8 1e8 1.00e-8 2.00e-5 1.97e-13 2.18
1e10 1e10 1.00e-10 1.00e-5 9.54e-16 2.03
Table 5: This table shows the influence and 1Exact scores for the red-cross point with the largest influence score when the black-dot point (see Figure 1 (left)) is placed at the $(x_{i},y_{i})$ position shown in the corresponding row of Table 4. In order to obtain the influence with respect to $\beta_{1}$, we let $e_{1}=(0,1)$, the standard basis vector corresponding to the x term in our linear regression setup. The Influence Score, $e_{1}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{j}(y_{j}-\hat{\beta}x_{j})$, is a product between the quantity in column 3 (which we call the leverage-like term, see Equation 13) and column 4 (the residuals). The 1Exact Score is the change in effect size that results from dropping the data point at $(x_{j},y_{j})$ and refitting OLS. The influence scores highlighted in yellow are those that are larger than the influence score of the black-dot point, leading AMIP and Greedy AMIP to misidentify the Most Influential Set of size 1, resulting in a Type 2 failure.
Red Cross:
$x_{j}$ $y_{j}$ $e_{1}^{\top}(\mathbf{X}^{\top}\mathbf{X})^{-1}x_{j}$ $(y_{j}-\hat{\beta}x_{j})$ Influence Score 1Exact Score
2.13 -0.09 2.02e-3 1.66 3.38e-3 3.40e-3
-0.72 -2.05 7.10e-5 -1.57 1.12e-4 1.12e-4
2.70 -4.85 7.25e-5 -7.66 5.55e-7 5.55e-7
2.70 -4.85 7.25e-8 -7.65 7.63e-9 7.64e-9
2.70 -4.85 9.97e-10 -7.65 7.65e-11 7.66e-11
2.70 -4.85 1.00e-14 -7.65 7.65e-13 7.66e-13